OpenStudy (anonymous):

How to solve x^2+24x+90=0 by completing the square? Can someone show me the steps to solve this?

OpenStudy (experimentx):

x^2+2*12x+(12)^2 - (12)^2+90 = 0

OpenStudy (anonymous):

I'm lost.

OpenStudy (anonymous):

\[x^2+2*12x+(12)^2 =54= ( 3 \sqrt 6)^2 \]

OpenStudy (kinggeorge):

You have an equation of the form \[x^2+bx+c=0\]To complete the square, calculate \[\left({b\over 2}\right)^2\]And then add and subtract this from your original equation. So now you have \[x^2+bx+{b^2 \over 4}-{b^2 \over 4}+c=0\]This factors as \[(x+{b \over 2})^2-{b^2 \over 4} +c=0\]

OpenStudy (anonymous):

(12)^2 - (12)^2 =0 so you added zero and nothing changes

OpenStudy (anonymous):

\[(x+12)^2 = (3 \sqrt 6)^2\]

OpenStudy (anonymous):

Noe you can finish t.

OpenStudy (kinggeorge):

By plugging in \(b=24\) and \(c=90\), you can solve\[(x+{b \over2})^2-{b^2 \over 4}+c=0\]For your question.

OpenStudy (anonymous):

is it x+12=36? and then you just solve it?

OpenStudy (anonymous):

I f \[ A^2 = B^2 \] tehn \[ A = \pm B \]

OpenStudy (anonymous):

I don't get it?

OpenStudy (anonymous):

\[ x+12 = \pm 3 \sqrt 6\]

OpenStudy (anonymous):

Oh ohkay. How would I go about solving that?

OpenStudy (anonymous):

There are two roots \[ x= -12 - 3 \sqrt 6, \quad x= -12 + 3 \sqrt 6, \]

OpenStudy (anonymous):

Thank you! :D

OpenStudy (anonymous):

As @KingGeorge described before, this how one proves the quadratic formula.